Please show me how to do in Excel or Google sheets (ie =normsinv, or =confidence, etc) TY!!

A.Giving a test to a group of students, the grades and gender are summarized below

A B C Total
Male 11 4 5 20
Female 3 12 14 29
Total 14 16 19 49

Let ππ represent the percentage of all male students who would receive a grade of C on this test. Use a 90% confidence interval to estimate ππ to three decimal places.

Enter your answer as a tri-linear inequality using decimals (not percents).___<pi<___

B. Out of 400 people sampled, 104 had kids. Based on this, construct a 95% confidence interval for the true population proportion of people with kids.
Give your answers as decimals, to three places
__ < p < __
Answer #1

Ans:

A)

sample proportion=5/20=0.25

90% confidence interval for p

=0.25+/-1.645*sqrt(0.25*(1-0.25)/20)

=0.25+/-0.159

=(0.091, 0.409)

0.091<p<0.409

B)sample proportion=104/400=0.260

95% confidence interval for p

=0.260+/-1.96*sqrt(0.260*(1-0.260)/400)

=0.260+/-0.043

=(0.217, 0.303)

0.217<p<0.303